The potential energy of a system of two particles

Help me
Answer #1

|F| = dU(r) / dr = -12 A (r ^-13)


It is pointed away from the particle
Answer #2

as F = kq1q2/r^2]


but enegry U = kq1q2/2

so Force = eenrgy*1/r

so Force F = U/r12    or A/r12^2
Answer #3

F =     \(-\frac{\partial U}{\partial r}\) =(12A)/r^13